Change page style: 

Gemini's Impact on Deep Impact

At Gemini the night of July 4th, 2005 was filled with fireworks, both earthly and celestial. In support of NASA’s Deep Impact (DI) mission two teams of scientists used the Gemini mid-infrared instruments MICHELLE and T-ReCS to acquire imaging and spectroscopic observations of Comet Tempel 1 throughout the impact event. At Gemini South a team led by David Osip (Carnegie Inst.) and James De Buizer (Gemini) used T-ReCS to image the impact and the resulting dust plume (Figure 1).

Announcement to NIRI Users in Semester 2001A and 2001B

Matt Mountain, Director

During our initial acceptance tests of NIRI on Gemini North, it was found that NIRI exhibited "flexure", or image motion instability problems that were far larger than expected and outside the original NIRI science requirements. The University of Hawaii's Institute for Astronomy (IfA), which is responsible for NIRI, has tried a number of measures to correct this problem. The image motion is caused by a poorly operating cryogenic bearing on the steering mirrors that select cameras within the NIRI cryostat.

Progress with NIRI during semester 2001B

30 Jan 2002

The 2001B semester has come to a close. Although no science observations could be completed, we have made progress towards our goal of starting regular science operations with NIRI.

NIRI Update

10 April 2002

NIRI Update

12 July 2002

Semester 2002A has now ended for NIRI. Our terrible luck with the weather continued through the last night of NIRI observations, and only a handful of the 41 assigned nights were useful for science observations. We did complete some observations for 15 queue programs. The data we have taken look quite good, although some were taken during poorer weather conditions than we would have liked. The data are now being distributed on CD ROM to the PIs. See the schedules pages for more information on a given program.

NIRI Update

12 September 2002

The NIRI maintenance work has now been completed and NIRI has been tested cold in the lab. All the mechanisms are moving and the new bushings are working well. In addition to the preventative maintenance, we also installed the new Shack-Hartmann roof prism in the on-instrument wavefront sensor optics. We look forward to using the 2x2 spot pattern to correct for fast tip-tilt and focus in the coming months. The NIRI science-channel camera optics were realigned and the pupil alignment is nearly identical to that before the shutdown.

NIRI Update

24 September 2002